OpenStudy (anonymous):

what is the value of the fourth term in a geometric sequence for which a 1 10 and r 0.5

OpenStudy (jhannybean):

First off, what is the equation for the geometric sequence?

OpenStudy (solomonzelman):

\(\large\color{black}{ a_1 =10}\) , \(\large\color{black}{ r =0.5}\) there is a form for any nth term: \(\large\color{black}{ a_{n}=a_1\times r^{n-1}}\)

OpenStudy (solomonzelman):

in your case, it is : \(\large\color{black}{ a_{4}=a_1\times r^{4-1}}\) \(\large\color{black}{ \Downarrow}\) \(\large\color{black}{ a_{4}=a_1\times r^{3}}\)

OpenStudy (solomonzelman):

you know the \(\large\color{black}{a_1}\) and the \(\large\color{black}{ r}\). If you have questions, ask.